# Unlocking the Power of Mixture-of-Experts for Task-Aware Time Series Analytics

### v2

@@ Appendix B Full Results
ETTh1
96 $\color[rgb]{0,0,1}\underline{0.355}$} $\color[rgb]{0,0,1}\underline{0.390}$} 0.345 0.373 0.370 0.396 0.373 0.399 0.386 0.405 0.372 0.392 0.372 0.401 0.377 0.397 0.411 0.435 0.389 0.412 0.379 0.403 0.379 0.419
@@
ETTh2
96 $\color[rgb]{0,0,1}\underline{0.272}$} 0.330 0.276 0.340 0.280 0.343 0.287 0.349 0.292 0.347 0.279 0.336 0.270 0.342 0.274 0.337 0.728 0.603 0.334 0.370 0.300 0.364 0.337 0.380
@@
Weather
96 $\color[rgb]{0,0,1}\underline{0.145}$} 0.183 0.151 0.203 0.151 0.202 0.147 0.199 0.157 0.207 0.148 0.195 0.147 0.198 0.150 0.200 0.143 0.210 0.168 0.214 0.170 0.230 0.229 0.298
@@
Electricity
96 0.131 $\color[rgb]{0,0,1}\underline{0.226}$} - - 0.135 0.231 0.132 0.227 0.134 0.230 0.135 0.222 0.153 0.256 0.143 0.247 0.134 0.231 0.169 0.271 0.140 0.237 0.191 0.305
@@
Traffic
96 0.361 $\color[rgb]{0,0,1}\underline{0.261}$} - - 0.388 0.269 0.391 0.277 0.363 0.265 0.392 0.271 0.369 0.256 0.370 0.262 0.526 0.288 0.595 0.312 0.395 0.275 0.593 0.365

### v3

# Task-Aware Mixture-of-Experts for Time Series Analysis

@@ Appendix B Full Results
ETTh1 96 $\color[rgb]{0,0,1}\underline{0.355}$} $\color[rgb]{0,0,1}\underline{0.390}$} 0.345 0.373 0.370 0.396 0.373 0.399 0.386 0.405 0.372 0.392 0.372 0.401 0.377 0.397 0.411 0.435 0.389 0.412 0.379 0.403 0.379 0.419
@@
ETTh2 96 $\color[rgb]{0,0,1}\underline{0.272}$} 0.330 0.276 0.340 0.280 0.343 0.287 0.349 0.292 0.347 0.279 0.336 0.270 0.342 0.274 0.337 0.728 0.603 0.334 0.370 0.300 0.364 0.337 0.380
@@
Weather 96 $\color[rgb]{0,0,1}\underline{0.145}$} 0.183 0.151 0.203 0.151 0.202 0.147 0.199 0.157 0.207 0.148 0.195 0.147 0.198 0.150 0.200 0.143 0.210 0.168 0.214 0.170 0.230 0.229 0.298
@@
Electricity 96 0.131 $\color[rgb]{0,0,1}\underline{0.226}$} - - 0.135 0.231 0.132 0.227 0.134 0.230 0.135 0.222 0.153 0.256 0.143 0.247 0.134 0.231 0.169 0.271 0.140 0.237 0.191 0.305
@@
Traffic 96 0.361 $\color[rgb]{0,0,1}\underline{0.261}$} - - 0.388 0.269 0.391 0.277 0.363 0.265 0.392 0.271 0.369 0.256 0.370 0.262 0.526 0.288 0.595 0.312 0.395 0.275 0.593 0.365